OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = x^2 - 6x + 9 ? a. (3, 0) b. (-3, 3) c. (4, 1) d. (1, 4)

hero (hero):

use x = -b over 2a to find the x-coordinate

OpenStudy (anonymous):

b

hero (hero):

What? Why is b the answer?

OpenStudy (anonymous):

because it looked good

OpenStudy (anonymous):

like a cheeseburger

hero (hero):

Looking good doesn't mean right, just like when you girls say looks ain't everything.

OpenStudy (anonymous):

looks are everything

hero (hero):

Not in this case it isn't

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Whats the answer? Wtf

hero (hero):

I gave you the formula to find the x coordinate

hero (hero):

Use it....

OpenStudy (anonymous):

yes use his formula

hero (hero):

Hint: It's not b

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

its b and a half?

hero (hero):

The answer isn't b anything

hero (hero):

Jasmine, you fail to properly apply my formula

OpenStudy (anonymous):

|b|-(0,3)?

hero (hero):

What the....? Where does that come from and what methods have you used?

OpenStudy (anonymous):

i made an algorithm

hero (hero):

You shouldn't have....lol

OpenStudy (anonymous):

i put my pencil on the paper and i draw circles and then i count the circles to see how many times i need to count through the choices and then i do -(0,3)

OpenStudy (anonymous):

the absolute value of that choice -(0,3)

hero (hero):

Your methods are questionable at the moment

OpenStudy (anonymous):

to tell you the truth i think this is the only time it works

hero (hero):

If something has only worked one time, It's probably not a valid method...

hero (hero):

Especially in math

OpenStudy (anonymous):

its valid for that one time

hero (hero):

Well, for this problem you use x = the opposite of b over 2a

OpenStudy (anonymous):

ok

hero (hero):

You can either use that or graph it to find x

OpenStudy (anonymous):

where does a and b come from

hero (hero):

a = 1 b = -6

OpenStudy (anonymous):

where is that

hero (hero):

The are coefficients of the first and second terms

OpenStudy (anonymous):

oh does that come from f(x)=x^2-6x+9

OpenStudy (anonymous):

what do we do with the 9?

hero (hero):

Yes

hero (hero):

We don't do anything with the 9. It's not part of the formula

OpenStudy (anonymous):

ok so we do b/2a -6/2(1)=-3(1)=-3

hero (hero):

x = -b/2a not b/2a

hero (hero):

-b means the opposite of b.

OpenStudy (anonymous):

oops i see a mistake of mine -3+1=-2 but i did it wrk it should be 3+1=4

OpenStudy (anonymous):

not wait its multiplication right so 3

hero (hero):

It's about time.....

hero (hero):

You finally got it

OpenStudy (anonymous):

but thats what i got doing my algorithm

OpenStudy (anonymous):

|b|-(0,3)

hero (hero):

I'd rather use -b/2a than some algorithm

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

gj hero